Subject: Diffusion Morning Report 12/20/12
From: Ted Berg <tberg@stanford.edu>
Date: Thu, 20 Dec 2012 07:07:33 -0800
Thu, 20 Dec 2012 07:07:33 -0800
Good Morning All,
*1**. New issues**: lab shutdown
* *2. Ongoing issues:*Tylan Poly rewriting recipes,** Tylannitride  
tested again waiting further instructions after alarms, *O2 in Silane?  
,**Epi dome coated-PM still in progress, ?Epi chamber 2 permit at 
county  DEH approval completed, *
***  3. Resolved:* **


Good Morning All,
      1. New issues: lab shutdown
  2. Ongoing issues:Tylan Poly rewriting recipes,  Tylannitride  tested again waiting further instructions after alarms, O2 in Silane?  , Epi dome coated-PM still in progress, ?Epi chamber 2 permit at county  DEH approval completed,
   3. Resolved: